Find the time tH it takes the projectile to reach its maximum height H. Express tH in terms of v0, θ, and g (the magnitude of the acceleration due to gravity).
Answer #1

at maximum height the verical component of velocity will becomes zero

in vertical direction

using kinematic equations

then vy = uy - g*tH

vy = 0 m/s at maximum height


uy is the initial component of vertical velocity = vo*sin(theta)

then 0 = vo*sin(theta) - g*tH

tH = vo*sin(theta)/g